OpenStudy (anonymous):

Find the exact value of the composition. arcos(cos pi/2)

OpenStudy (jhannybean):

These functions undo eachother. This is the definition of the inverse function. \[\large \cos^{-1}(\cos(x))=x\]and\[\large \cos(\cos^{-1}(x))=x\] For example, say you wanted to find f(g(x)) where f(x)=x+2 and g(x)=x-2 f(g(x))= (x-2) + 2 = x Trig functions work the same way.